5) For each of the following recurrences state whether the Master theorem can be applied to solve the recurrence or not. If the Master theorem can be used, then use it to determine running time for the recurrence. If the Master theorem cannot be applied, then specify the reason (you dont need to solve the recurrence). a) T(n) = 4T(n/3)+n2
Answer #1

Yes, Master theorem cn be applied to

T(n) = 4T(n/3) + n^2


T(n) = aT(n/b) + f(n) where a >= 1 and b > 1

Here, a = 4, b = 3, f(n) = n^2